OpenStudy (anonymous):

Which of the following is the equation of g(x) if the parent function is f(x) = x^2 and a=1/4 , h = -2, and k = -10?

OpenStudy (anonymous):

these are the multiple choice answers

OpenStudy (anonymous):

@Abhisar @Destinymasha @abb0t @uri @Isaiah.Feynman @confluxepic anyone?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@myininaya @ganeshie8 @saifoo.khan

myininaya (myininaya):

you are to replace letters in an equation

myininaya (myininaya):

what did you get when you replaced the letters given in your problem into the equation

myininaya (myininaya):

the equation you most likely are to use is : \[g(x)=a(x-h)^2+k\]

myininaya (myininaya):

you replaced the a with what it says to replace it with which is 1/4 replace h with -2 replace k with -10 I know to do this because of the equalities given in your problem

OpenStudy (anonymous):

so g(x) = a (x- -2)^2 + -10

myininaya (myininaya):

why not replace the a also?

OpenStudy (anonymous):

oh, missed that one

OpenStudy (anonymous):

the answer would be d then, right? g(x)= 1/4(x+2)^2 -10

OpenStudy (anonymous):

@myininaya

myininaya (myininaya):

i'm not looking at your choice selection but yeah you replaced a with 1/4 and replaced h with -2 and k with -10

myininaya (myininaya):

I don't see anything wrong with your replacements ...

OpenStudy (anonymous):

Ok and Thank You, I see you are very busy and truly understand. =)

OpenStudy (anonymous):

thanks(: